Beth has hamsters at her house. The number of hamsters is equal to the sum of the number of birds at kristens house and the number of dogs at wendys house. John has the same number of cats as wendy has of dogs. If these four children have a total of 12 pets, how many pets are at each house?
Answer by stanbon(75887)

Beth has hamsters at her house. The number of hamsters is equal to the sum of the number of birds at kristens house and the number of dogs at wendys house. John has the same number of cats as wendy has of dogs. If these four children have a total of 12 pets, how many pets are at each house?
-----------
Equations:
h = b + d
c = d
h + b + c + d = 12
------
Substitute and solve for "d
b + d + +d + d = 12
------
b + 3d = 12
-----
Note: Since d is a whole number, d could be 1 or 2 or 3
--
If d = 1, b = 9
If d = 2, b = 6
If d = 3, b = 3
----------------
Cheers,
Stan H.
================